When a 3.00 g

3.00 g sample of KCN

KCN is dissolved in water in a calorimeter that has a total heat capacity of 1.23 kJ⋅K

−1

,

1.23 kJ⋅K−1, the temperature decreases by 0.440 K.

0.440 K. Calculate the molar heat of solution of KCN.


1
Expert's answer
2021-02-09T04:43:37-0500

Q=nC∆T.

n=m/Mr = 3/65 = 0,0461538462 moles.

C= Q/n∆T= 1.23/0,0461538462 x 0.44 = 49,2424241